OpenStudy (anonymous):

A certain fungus grows in a circular shape. Its diameter in inches after t weeks is given below. 6- 50/t^2+10 What is the area covered by the fungus when t = 0? What area does it cover at the end of 5 weeks?

OpenStudy (anonymous):

The area covered is 0 square inches at t = 0 and 5.2π square inches at t = 5. The area covered is square inches at t = 0 and 4.2π square inches at t = 5. The area covered is square inches at t = 0 and 5.2π square inches at t = 5. The area covered is square inches at t = 0 and 10.4π square inches at t = 5. The area covered is square inches at t = 0 and 5.2π square inches at t = 5.

OpenStudy (anonymous):

are the answers

OpenStudy (anonymous):

the blanks are: pi/4, pi/4, pi/6, and pi/2 (in the answers above) ...sorry copied and pasted

OpenStudy (amistre64):

Area = pi (d/2)^2

OpenStudy (amistre64):

depending on how yout equation is actually spose to look: id say its something like this: Area = pi*[(6- 50/t^2+10)/2]^2

OpenStudy (anonymous):

i am confused

OpenStudy (amistre64):

all i did was take the function they gave you for the diameter and plugged it straight into the Area function of a circle

OpenStudy (amistre64):

i got know idea what you equation is spose to look like, but thats on your end. plug in the t values and itll just sipt out an answer

OpenStudy (anonymous):

i know i know but the answers they are giving me are confusing me

OpenStudy (amistre64):

what do you get for d when t = 0?

OpenStudy (anonymous):

1

OpenStudy (amistre64):

yeah, 1 is good; now use that in the circles area formula A = (d/2)^2 pi A = 1/4 pi

OpenStudy (amistre64):

3.14/4 = .78 something or another right?

OpenStudy (amistre64):

or .25 pi

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

what you get for t=5 in the diamter equation?

OpenStudy (anonymous):

4.6 or something like that

OpenStudy (anonymous):

i rounded

OpenStudy (amistre64):

k, and 4.6/2 = 2.3 (2.3)^2 = 5.29 or so, so at t=5 the area = 5.29 pi or so

OpenStudy (amistre64):

its either the middle or the last, since yout t=0 stuff got eaten

OpenStudy (anonymous):

yeah that is right.

OpenStudy (anonymous):

how did u get 1/4 for the first thing again?

OpenStudy (anonymous):

nevermind

OpenStudy (amistre64):

did you figure it out?